undeadknight26 (undeadknight26):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

undeadknight26 (undeadknight26):

How to solve the rest of 2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y

undeadknight26 (undeadknight26):

@e.mccormick @E.ali @ehuman @agent0smith @UsArmy3947

OpenStudy (agent0smith):

What do you mean the rest? it looks like you're done.

undeadknight26 (undeadknight26):

oh ok...

OpenStudy (agent0smith):

For y = mx + b, m is the slope, b is the y intercept. They should be easy to identify now.

undeadknight26 (undeadknight26):

so the slope is .66 and the x is 400? what would the y be though?

undeadknight26 (undeadknight26):

no y would be 400?

undeadknight26 (undeadknight26):

@agent0smith please come back!

OpenStudy (anonymous):

I just did this assignment yesterday. What you have so far is, if put into y = mx + b form: y = -2/3x + 400 If b = 400, and b is the y-intercept, then your y-intercept should be at the ordered pair (0,400) and your slope is -2/3. Is that what you were confused about?

undeadknight26 (undeadknight26):

so i have no x?

OpenStudy (anonymous):

For this question, you don't necessarily need to solve for x, it is only asking you to "identify the slope and y-intercept of the equation"

OpenStudy (agent0smith):

For y = mx + b, m is the slope, b is the y intercept. That's all you're asked for.

OpenStudy (anonymous):

You only have to say that slope m = -2/3 and y-intercept b = 400, I promise :3

undeadknight26 (undeadknight26):

I trust you!